OpenStudy (anonymous):

6. A photon of light has a wavelength of about 600 nm (1nm = 1 x 10-9 m). What is the energy of this photon of light? Show all calculations leading to an answer.

OpenStudy (anonymous):

\[E = \frac{ hc }{ \lambda }\] \[h = Plank's~constant=6.626 \times 10^{-34}~J-s\] \[c=speed~of~light = 3.0 \times 10^8~m/s\] \[\lambda=wavelength~[meters]\]

OpenStudy (anonymous):

First of all.. "photon of light" (funny thing..) Next.. Upvote for @peachpi .. u are right..